If the two stage rocket of a weights 100 kg and 10 kg separately and contains 800 kg and 90 kg fuel


separately, what is the final velocity that can be achieved with an exhaust velocity of 1.5 km/sec. (Given


log10 =2.3 and log102 =0.3 )

1
Expert's answer
2022-04-05T19:19:14-0400

"m_1=100,~M_1=800,~m_2=10,~M_2=90,"

"v=v_0\ln\frac{m_1+M_2+m_2+M_2}{m_1+m_2+M_2}\ln\frac{m_2+M_2}{m_2}=5. 56~\frac{km}s."